According to this relation the value of Y when X = 5 is? Then teacher gives a table X= 0, 1, 2, 3, 4 than Y=0, 3, 12, 27, 48. We have to add 5 for X and what for Y? please help. We have to show work and answer a multiply choice question which is A=67, B=70, C=73, D=75
Thank you

Answer by rothauserc(4718)

We have the sequence
:
0, 3, 12, 27, 48
;
we use the differences approach - we look for a level where the differences are the same
:
first level differences are 3, 9, 15, 21
second level differences are 6, 6, 6
:
we know the function is of the form y = k * x^2, where k is a constant
:
by inspection, the function is
:
y = 3 * x^2
:
therefore, for x = 5
:
y = 3 * 5^2 = 75
: